# FSC: Few-point Shape Completion

### v1

@@ Abstract
While previous studies have demonstrated successful 3D object shape completion with a sufficient number of points, they often fail in scenarios when a few points, e.g. tens of points, are observed. Surprisingly, via entropy analysis, we find that even a few points, e.g. 64 points, could retain substantial information to help recover the 3D shape of the object. To address the challenge of shape completion with very sparse point clouds, we then propose Few-point Shape Completion (FSC) model, which contains a novel dual-branch feature extractor for handling extremely sparse inputs, coupled with an extensive branch for maximal point utilization with a saliency branch for dynamic importance assignment. This model is further bolstered by a two-stage revision network that refines both the extracted features and the decoder output, enhancing the detail and authenticity of the completed point cloud. Our experiments demonstrate the feasibility of recovering 3D shapes from a few points. The proposed Few-point Shape Completion (FSC) model outperforms previous methods on both few-point inputs and many-point inputs, and shows good generalizability to different object categories.
@@ 1 Introduction
Despite the progress in point cloud-based shape completion highlighted in works like [57, 1, 31, 20, 35, 32, 52, 67], few have tackled the challenge of completing a point cloud with few input points. Previous works like [57, 51] typically utilize a few thousand points as input and demonstrate competent completion capabilities with this volume of input. However, these works do not investigate the potential for point cloud completion when the number of input points is significantly reduced, which scenario might arise in many practical applications. They do not explore the extent of the information contained in a few-point input or attempt to leverage this information to enhance point cloud completion. Previous studies like PCN [57] use feature encoders, often inspired by PointNet [24] or PointNet++[26], for capturing input shape features. However, with a limited number of points, where each point is more crucial, the feature extraction must effectively use all points and prioritize their importance for diverse shape information. This ensures the feature quality from few-point inputs is on par with many-point inputs. Conversely, voxel-based methods like GRNet [51] may succeed in global shape recovery, but they face challenges in creating dense representations from a few points. Fig. 1 shows 64 points input results in inaccurate global shapes with PCN, while voxel-based approaches GRNet yield sparse outcomes.
@@ 2.1 Point-cloud-based Shape Completion
Point cloud completion [22, 21, 61, 59, 34, 65, 46, 58, 38, 3, 43, 6, 36, 50, 33] is a vital task in 3D computer vision with a multitude of applications, such as robot perception, autonomous driving, and augmented reality. Voxel-based and point-based methods are the two primary categories of point cloud completion approaches. The voxel-based method converts the input point cloud into voxels and subsequently generates complementary results in both voxel and point cloud representations. GRNet [51] exemplifies a voxel-based method that leverages 3D CNNs to extract voxel features and multilayer perceptrons (MLPs) to produce complementary results in point cloud representation. Nonetheless, voxel-based methods exhibit a lack of robustness when confronted with a limited number of input points, leading to missing voxel neighborhoods and, consequently, the generation of results containing holes.
In contrast, point-based methods directly perform point cloud completion on the input points. PCN [57], a trailblazing point-based method, segregates the completion process into an encoder and a decoder. The encoder utilizes PointNet to extract global features, while the decoder employs fully connected and FoldingNet networks to generate coarse and fine point clouds, respectively. Given that the completion process negates the need for neighborhood calculation, point-based methods are more resilient to variations in the number of input points. However, their completion outcomes are often suboptimal, prompting the advent of various enhancement methods such as PF-Net [14], TopNet [25], and SA-Net [62]. These methods concentrate on procuring more representative point clouds with distinct geometric features and progressively amplifying the resolution of the generated point cloud.
@@ 2.2 Point Cloud Analysis
Point cloud analysis encompasses a range of tasks including segmentation, object detection, registration, completion, and surface reconstruction, with deep learning-based methods, such as [24, 40, 15, 54, 13, 42, 19, 17, 18], being widely implemented. In these works, point clouds are typically encoded either by leveraging their connections with neighboring points, as seen in methods like PointNet++[26], DGCNN[40], and Point Transformer [63], or by employing simple networks such as MLP, as in the encoder of PCN [57]. While these encoding designs have proven to be effective in previous point cloud completion tasks, they tend to fall short when dealing with inputs comprising a limited number of points. Given this constraint of limited input information, our point cloud encoder must be significantly more effective than those employed in previous point cloud analysis approaches.
@@ 3 Few-point Information Analysis
[…]
@@ 4.4 Training Losses
$\small\begin{split}&{d_{1}}\left({{Y_{{\rm{coarse}}}},{{\tilde{Y}}_{{\rm{gt}}}}}\right)={\rm{EMD}}\left({{Y_{{\rm{coarse}}}},{{\tilde{Y}}_{{\rm{gt}}}}}\right)\\ &=\mathop{\min}\limits_{\phi:{Y_{{\rm{coarse}}}}\to{{\tilde{Y}}_{{\rm{gt}}}}}\frac{1}{{\left|{{Y_{{\rm{coarse}}}}}\right|}}\sum\limits_{p\in{Y_{{\rm{coarse}}}}}{{{\left\|{p-\phi\left(p\right)}\right\|}_{2}}}.\end{split}$ (7)
@@
$\footnotesize\begin{split}&{d_{2}}\left({{Y_{{\rm{detail}}}},{Y_{{\rm{gt}}}}}\right)={\rm{CD}}\left({{Y_{{\rm{detail}}}},{Y_{{\rm{gt}}}}}\right)\\ &=\frac{1}{{\left|{{Y_{{\rm{detail}}}}}\right|}}\sum\limits_{p\in{Y_{{\rm{detail}}}}}{\mathop{\min}\limits_{q\in{Y_{{\rm{gt}}}}}{{\left\|{p-q}\right\|}_{2}}}{\rm{+}}\frac{1}{{\left|{{Y_{{\rm{gt}}}}}\right|}}\sum\limits_{q\in{Y_{{\rm{gt}}}}}{\mathop{\min}\limits_{p\in{Y_{{\rm{detail}}}}}{{\left\|{q-p}\right\|}_{2}}}.\end{split}$ (8)
@@
Table 1: 64 points completion accuarry on the 8 “seen categories” of ShapeNet. We can see that our method outperforms previous methods on average and in most categories. Avg CD-${\ell_{1}}$ the lower the better. Bold number means the best.
Table 2: 2048 points completion accuarry on the 8 “seen categories” of ShapeNet. We can see that our method outperforms previous methods in most categories and the average results.
Avg CD-${\ell_{1}}\downarrow$ Airplane Cabinet Car Chair Lamp Sofa Table Vessel 3D-EPN [5] 20.15 13.16 21.80 20.31 18.81 25.75 21.09 21.712 18.54 FoldingNet [53] 14.31 9.49 15.80 12.61 15.55 16.41 15.97 13.65 14.99 AtlasNet [9] 10.85 6.37 11.94 10.10 12.06 12.37 12.99 10.33 10.61 CRN [37] 11.98 6.44 15.09 13.75 12.37 11.38 14.70 11.78 10.33 MSN [16] 10.00 5.60 11.90 10.30 10.20 10.70 11.60 9.60 9.90 PCN [57] 9.64 5.50 10.63 8.70 11.00 11.34 11.68 8.59 9.67 TopNet [25] 9.89 6.24 11.63 9.83 11.50 9.37 12.35 9.36 8.85 GRNet [51] 8.83 6.45 10.37 9.45 9.41 7.96 10.51 8.44 8.04 PMP-Net [44] 8.66 5.50 11.10 9.62 9.47 6.89 10.74 8.77 7.19 PoinTr [55] 8.38 4.75 10.47 8.68 9.39 7.75 10.93 7.78 7.29 SCRN [39] 8.29 4.80 9.94 9.31 8.78 8.66 9.74 7.20 7.91 NSFA [60] 8.06 4.76 10.18 8.63 8.53 7.03 10.53 7.35 7.48 SnowFlake [49] 7.21 4.29 9.16 8.08 7.89 6.07 9.23 6.55 6.40 Ours 7.02 4.07 9.12 8.10 7.21 5.88 9.30 6.26 6.25
@@
Table 2: 2048 points completion accuarry on the 8 “seen categories” of ShapeNet. We can see that our method outperforms previous methods in most categories and the average results.
@@ 5.1 Data Generation and Datasets
ShapeNet data generation. We train and evaluate our method on ShapeNet [2]. It is a synthetic CAD dataset comprising 30,974 complete and detailed 3D point clouds. There are in total 16 categories of point clouds in ShapeNet. We use 8 of them for training and testing (“seen categories”), and use the other 8 categories (“unseen categories”) for the generalizability test. To simulate realistic input point clouds, we sample input point clouds from the mesh surface. Specifically, we first uniformly sample 16,384 points as Ground truth. Then, we generated partial point clouds with 2048 points by back-projecting 2.5D depth images into 3D. Subsequently, we downsampled these to create low-resolution partial point clouds with varying densities: 1024, 512, 256, 128, and 64 points. The entire dataset of 30,974 3D point clouds was divided into training (28,974 point clouds), validation (800 point clouds), and testing sets (1,200 point clouds). We use the same (CD-${\ell_{1}}$) as in PCN [57] to evaluate our completion accuracy on both datasets. CD-${\ell_{1}}$ is the Chamfer Distance using L1-norm, which is defined as
@@ 5.2 Shannon Information Analysis
Figure 6: Visualized results for 64 input points. The top four rows are the results of categories used during training (“seen” categories). The bottom four rows are the results of categories not seen during training (“unseen” categories). As we can see, our method has better visualization results than previous methods.
@@ 5.3 Results
Few-point results on ShapeNet. We train and evaluate our method on 8 categories of point clouds on ShapeNet. Tab. 1 presents the per-class completion results on ShapeNet for scenarios where the input point cloud contains only 64 points. In all eight “seen categories”, our method outperforms others, achieving the highest accuracy. We observe a notable improvement over the previous method. This demonstrates that our approach has a distinct advantage in completion accuracy, especially in cases of few-point point cloud input. We also compare the performance on ShapeNet when the input point number varies from 2048 to 16 in Fig. 4. We can observe that our FSC outperforms previous approaches when the number of input points changes. Besides, when the number input point is less than 64, the performance of all methods drops sharply.

### v5

@@ Abstract
While previous studies have demonstrated successful 3D object shape completion with a sufficient number of points, they often fail in scenarios when a few points, e.g. tens of points, are observed. Surprisingly, via entropy analysis, we find that even a few points, e.g. 64 points, could retain substantial information to help recover the 3D shape of the object. To address the challenge of shape completion with very sparse point clouds, we then propose Few-point Shape Completion (FSC) model, which contains a novel dual-branch feature extractor for handling extremely sparse inputs, coupled with an extensive branch for maximal point utilization with a saliency branch for dynamic importance assignment. This model is further bolstered by a two-stage revision network that refines both the extracted features and the decoder output, enhancing the detail and authenticity of the completed point cloud. Our experiments demonstrate the feasibility of recovering 3D shapes from a few points. The proposed Few-point Shape Completion (FSC) model outperforms previous methods on both few-point inputs and many-point inputs, and shows good generalizability to different object categories. Code is available at https://github.com/xianzuwu/FSC.
@@ 1 Introduction
Despite the progress in point cloud-based shape completion highlighted in works like [59, 1, 67], few have tackled the challenge of completing a point cloud with few input points. Previous works like [59, 54] typically utilize a few thousand points as input and demonstrate competent completion capabilities with this volume of input. However, these works do not investigate the potential for point cloud completion when the number of input points is significantly reduced, which scenario might arise in many practical applications. They do not explore the extent of the information contained in a few-point input or attempt to leverage this information to enhance point cloud completion. Previous studies like PCN [59] use feature encoders, often inspired by PointNet [30] or PointNet++[31], for capturing input shape features. However, with a limited number of points, where each point is more crucial, the feature extraction must effectively use all points and prioritize their importance for diverse shape information. This ensures the feature quality from few-point inputs is on par with many-point inputs. Conversely, voxel-based methods like GRNet [54] may succeed in global shape recovery, but they face challenges in creating dense representations from a few points. Fig. 1 shows 64 points input results in inaccurate global shapes with PCN, while voxel-based approaches GRNet yield sparse outcomes.
@@ 2.1 Point-cloud-based Shape Completion
Point cloud completion [27, 26, 62, 61, 39, 65, 60, 42, 3, 47, 9, 40, 53, 38] is a vital task in 3D computer vision with a multitude of applications, such as robot perception, autonomous driving, and augmented reality. Voxel-based and point-based methods are the two primary categories of point cloud completion approaches. The voxel-based method converts the input point cloud into voxels and subsequently generates complementary results in both voxel and point cloud representations. GRNet [54] exemplifies a voxel-based method that leverages 3D CNNs to extract voxel features and multilayer perceptrons (MLPs) to produce complementary results in point cloud representation. Nonetheless, voxel-based methods exhibit a lack of robustness when confronted with a limited number of input points, leading to missing voxel neighborhoods and, consequently, the generation of results containing holes.
In contrast, point-based methods directly perform point cloud completion on the input points. PCN [59], a trailblazing point-based method, segregates the completion process into an encoder and a decoder. The encoder utilizes PointNet to extract global features, while the decoder employs fully connected and FoldingNet networks to generate coarse and fine point clouds, respectively. Given that the completion process negates the need for neighborhood calculation, point-based methods are more resilient to variations in the number of input points. However, their completion outcomes are often suboptimal, prompting the advent of various enhancement methods such as PF-Net [17], PointWOLF [18], RSMix [19],TopNet [37] and SA-Net [46] . These methods concentrate on procuring more representative point clouds with distinct geometric features and progressively amplifying the resolution of the generated point cloud.
@@ 2.2 Point Cloud Analysis
Point cloud analysis encompasses a range of tasks including segmentation, object detection, registration, completion, and surface reconstruction, with deep learning-based methods, such as [30, 44, 56, 24, 22, 23, 7, 25, 32, 4, 29, 20, 33, 12, 5], being widely implemented. In these works, point clouds are typically encoded either by leveraging their connections with neighboring points, as seen in methods like PointNet++[31], DGCNN[44], and Point Transformer [63], or by employing simple networks such as MLP, as in the encoder of PCN [59]. While these encoding designs have proven to be effective in previous point cloud completion tasks, they tend to fall short when dealing with inputs comprising a limited number of points. Given this constraint of limited input information, our point cloud encoder must be significantly more effective than those employed in previous point cloud analysis approaches.
@@ 3 Few-point Information Analysis
$S=-\sum_{i}\textit{FPFH}_{i}\log(\textit{FPFH}_{i}),$ (1)
where $\textit{FPFH}_{i}$ is the $i$th bin of normalized FPFH.
@@ 4.4 Training Losses
$\begin{split}&{d_{1}}\left({{Y_{{\rm{coarse}}}},{{\tilde{Y}}_{{\rm{gt}}}}}\right)={\rm{EMD}}\left({{Y_{{\rm{coarse}}}},{{\tilde{Y}}_{{\rm{gt}}}}}\right)\\ &=\mathop{\min}\limits_{\phi:{Y_{{\rm{coarse}}}}\to{{\tilde{Y}}_{{\rm{gt}}}}}\frac{1}{{\left|{{Y_{{\rm{coarse}}}}}\right|}}\sum\limits_{p\in{Y_{{\rm{coarse}}}}}{{{\left\|{p-\phi\left(p\right)}\right\|}_{2}}}.\end{split}$ (7)
@@
$\begin{split}&{d_{2}}\left({{Y_{{\rm{detail}}}},{Y_{{\rm{gt}}}}}\right)={\rm{CD}}\left({{Y_{{\rm{detail}}}},{Y_{{\rm{gt}}}}}\right)\\ &=\frac{1}{{\left|{{Y_{{\rm{detail}}}}}\right|}}\sum\limits_{p\in{Y_{{\rm{detail}}}}}{\mathop{\min}\limits_{q\in{Y_{{\rm{gt}}}}}{{\left\|{p-q}\right\|}_{2}}}{\rm{+}}\frac{1}{{\left|{{Y_{{\rm{gt}}}}}\right|}}\sum\limits_{q\in{Y_{{\rm{gt}}}}}{\mathop{\min}\limits_{p\in{Y_{{\rm{detail}}}}}{{\left\|{q-p}\right\|}_{2}}}.\end{split}$ (8)
@@
Table 1: 64 points completion accuarry on the 8 “seen categories" of ShapeNet. We can see that our method outperforms previous methods on average and in most categories. Avg CD-${\ell_{1}}$ the lower the better. Bold number means the best.
Avg CD-${\ell_{1}}\downarrow$ Airplane Cabinet Car Chair Lamp Sofa Table Vessel
3D-EPN [8] 20.15 13.16 21.80 20.31 18.81 25.75 21.09 21.712 18.54
FoldingNet [55] 14.31 9.49 15.80 12.61 15.55 16.41 15.97 13.65 14.99
AtlasNet [13] 10.85 6.37 11.94 10.10 12.06 12.37 12.99 10.33 10.61
CRN [41] 11.98 6.44 15.09 13.75 12.37 11.38 14.70 11.78 10.33
MSN [21] 10.00 5.60 11.90 10.30 10.20 10.70 11.60 9.60 9.90
PCN [59] 9.64 5.50 10.63 8.70 11.00 11.34 11.68 8.59 9.67
TopNet [37] 9.89 6.24 11.63 9.83 11.50 9.37 12.35 9.36 8.85
GRNet [54] 8.83 6.45 10.37 9.45 9.41 7.96 10.51 8.44 8.04
PMP-Net [48] 8.66 5.50 11.10 9.62 9.47 6.89 10.74 8.77 7.19
PoinTr [57] 8.38 4.75 10.47 8.68 9.39 7.75 10.93 7.78 7.29
SCRN [43] 8.29 4.80 9.94 9.31 8.78 8.66 9.74 7.20 7.91
NSFA [61] 8.06 4.76 10.18 8.63 8.53 7.03 10.53 7.35 7.48
SnowFlake [52] 7.21 4.29 9.16 8.08 7.89 6.07 9.23 6.55 6.40
Ours 7.02 4.07 9.12 8.10 7.21 5.88 9.30 6.26 6.25
Table 2: 2048 points completion accuarry on the 8 “seen categories" of ShapeNet. We can see that our method outperforms previous methods in most categories and the average results.
@@ 5.1 Data Generation and Datasets
ShapeNet data generation. We train and evaluate our method on ShapeNet [2]. It is a synthetic CAD dataset comprising 30,974 complete and detailed 3D point clouds. There are in total 16 categories of point clouds in ShapeNet. We use 8 of them for training and testing (“seen categories"), and use the other 8 categories (“unseen categories") for the generalizability test. To simulate realistic input point clouds, we sample input point clouds from the mesh surface. Specifically, we first uniformly sample 16,384 points as Ground truth. Then, we generated partial point clouds with 2048 points by back-projecting 2.5D depth images into 3D. Subsequently, we downsampled these to create low-resolution partial point clouds with varying densities: 1024, 512, 256, 128, and 64 points. The entire dataset of 30,974 3D point clouds was divided into training (28,974 point clouds), validation (800 point clouds), and testing sets (1,200 point clouds). We use the same (CD-${\ell_{1}}$) as in PCN [59] to evaluate our completion accuracy on both datasets. CD-${\ell_{1}}$ is the Chamfer Distance using L1-norm, which is defined as
@@ 5.2 Shannon Information Analysis
Figure 6: Visualized results for 64 input points. The top four rows are the results of categories used during training (“seen" categories). The bottom four rows are the results of categories not seen during training (“unseen" categories). As we can see, our method has better visualization results than previous methods.
@@ 5.3 Results
Few-point results on ShapeNet. We train and evaluate our method on 8 categories of point clouds on ShapeNet. Tab. 1 presents the per-class completion results on ShapeNet for scenarios where the input point cloud contains only 64 points. In all eight “seen categories", our method outperforms others, achieving the highest accuracy. We observe a notable improvement over the previous method. This demonstrates that our approach has a distinct advantage in completion accuracy, especially in cases of few-point point cloud input. We also compare the performance on ShapeNet when the input point number varies from 2048 to 16 in Fig. 4. We can observe that our FSC outperforms previous approaches when the number of input points changes. Besides, when the number input point is less than 64, the performance of all methods drops sharply.